OpenStudy (prettygirl_shynice):

.You are raping boxes for fund raiser you have 80 larger boxes. You need 2 3/4 feet of ribbon for each box each roll of ribbon has 50 feet of ribbon. estimates the number feet of ribbon you will need find the exact number of feet you will need how many rolls will you need? plz dont answer just explain will medeal+fan

OpenStudy (boldjon):

You have 80 large boxes. You require 2 3/4 feet of ribbon for each box. Total of ribbon required = 80 x 2 3/4 feet = 220 feet of ribbon. Each roll of ribbon is 50 feet long 220 / 50 = 4.4 rolls of ribbon required. You need 5 rolls of ribbon.

OpenStudy (boldjon):

hope that helped

OpenStudy (prettygirl_shynice):

okay may explain

OpenStudy (purple_pink):

by the way you missed the w in wraping XD

OpenStudy (prettygirl_shynice):

opps ssryr

OpenStudy (prettygirl_shynice):

may u explain thge steps plz

OpenStudy (purple_pink):

ummmm

OpenStudy (prettygirl_shynice):

ok

OpenStudy (purple_pink):

You have 80 large boxes to wrap. You will be needing 2 3/4 feet of ribbon for each box. So then the total amount of ribbon required = 80 x 2 3/4 feet = 220 feet of ribbon. Each roll of ribbon is 50 feet long, so you divide 220 / 50 = 4.4 rolls of ribbon required. You need 5 rolls of ribbon.

OpenStudy (purple_pink):

did that help?

OpenStudy (purple_pink):

@Prettygirl_shynice

OpenStudy (prettygirl_shynice):

yea

OpenStudy (prettygirl_shynice):

oso do i explain like that

OpenStudy (purple_pink):

yup

OpenStudy (prettygirl_shynice):

thx